OpenStudy (hellomynameiskam):

Solve the equation: x+ 7/8=4 A.4 7/8 B. x= 3 1/8 C.x= - 3 1/8 D. - 4 7/8

OpenStudy (kendricklamar2014):

1. Subtract 7/8 from both sides: \[x + \frac{ 7 }{ 8 } - \frac{ 7 }{ 8 } = 4 - \frac{ 7 }{ 8 }\] \[x = \frac{ 25 }{ 8 }\]

OpenStudy (kendricklamar2014):

\[\frac{ 25 }{ 8 } = 3.125....\] \[3\frac{ 1 }{ 8 } = 3.125\] \[Answer: B\]

OpenStudy (hellomynameiskam):

thanks. Can u help me with 4 more?

OpenStudy (kendricklamar2014):

Sure

OpenStudy (kendricklamar2014):

OS Rule: One Question per post. If you have more question, just make another post...